40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 6 05:44:14 2009, Yoichi, Update, Locking, Locking distracted by the QPD whitenning problem again 
    Reply  Fri Mar 6 15:23:39 2009, Yoichi, Update, Locking, Locking distracted by the QPD whitenning problem again lockLoss3.pdflockLoss3-zoom.pdf
Message ID: 1364     Entry time: Fri Mar 6 05:44:14 2009     Reply to this: 1365
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking distracted by the QPD whitenning problem again 
Tonight, I was able to ramp up the arm power to around 20. Then the DARM loop started to oscillate and the IFO lost lock in a few seconds.
I repeated this several times, then realized that the transmission QPDs were not working properly again due to the well known sticky slider problem.
I should have run slider_twiddle script. Since the DARM RF signal is normalized by the sqrt(TRX+TRY), it is reasonable that the DARM loop got unstable.

The fact that I was able to go up to arm power = 20 means there is nothing saturating below this power level.
ELOG V3.1.3-